OpenStudy (anonymous):

An arrow is shot straight up at 50m/s. How long will it take before it hits the ground (assuming it is shot from ground level)?

OpenStudy (anonymous):

Delta velocity = accelration*time. (The accelration is constant=g) Therefore (-50)-50 =(-10)*t T=10 s

OpenStudy (anonymous):

when does it reach its max height, when final velocity is zero. we know acceleration is -9.8 m/s/s use vf=vi+a*t 0=50-9.8*t t = 5.10 seconds. now, we know that the time it takes to go up, must equal to the time it takes to fall down. as the same force is acting on it so t = 5.1*2 =10.2